Find the consecutive even numbers whose sum is 108.





Found 3 solutions by Fombitz, lefty4ever26, stanbon:
Answer by Fombitz(32388)

How many of them?

Answer by lefty4ever26(59)

You didn't say how many consecutive even numbers you needed...but the formula would still work the same. Your first even number would be x. so the next one would have to be x +2. and the next would be x+4. SO
x+(x+2)+(x+4)=108
3x+6=108
3x=102
x=34
SO
your 'three' numbers would be 34, 36, 38.
If you needed more than three numbers you would add another even digit by x+6,x+8 etc.

Answer by stanbon(75887)

Find the consecutive even numbers whose sum is 108.
-----
1st: 2x
2nd: 2x+2
-----
Equation:
2x + 2x+2 = 108
4x = 106
x = 26.5
----
1st: 2x = 53
2nd: 2x+2 = 55
-----
There are no two consecutive EVEN numbers whose sum is 108.
=============
Cheers,
Stan H.